# Gravastar on the brane with a timelike extra dimension

### v1

@@ 5.1 Surface Redshift
$Z_{s}$ $=$ $-1+\frac{1}{\sqrt{g_{\it tt}}}=-1+{\frac{1}{\sqrt{{C}^{2}{{\rm e}^{Br^{2}}}}}}.$ (5.1)
Figure 3: Variation of the Surface redshift of the shell with respect to $r$.
@@
The radial profile of the surface redshift $Z_{s}$ from the inner to the outer boundary of the gravastar is illustrated in Figure 3. For any self-gravitating compact object, it is well known that the surface redshift must satisfy $Z_{s}\leq 2$ [63, 64, 65] in the absence of a cosmological constant. In our analysis within the Shtanov–Sahni braneworld framework, we do not include a $\Lambda$ term, as the effective four-dimensional cosmological constant vanishes due to a fine-tuned cancellation between the negative brane tension and the positive cosmological constant of the five-dimensional de Sitter bulk ($dS_{5}$).
@@
Using the numerically determined model parameters obtained from the matching and boundary conditions, the surface redshift reaches its maximum at the inner boundary of the gravastar and decreases smoothly across the shell, attaining a finite nonzero value at the outer boundary. The monotonic decrease of $Z_{s}$ throughout the shell indicates a well-behaved gravitational potential, while its magnitude remains comfortably below the theoretical upper bound. This behavior not only confirms the regularity of the shell but also provides strong evidence for the physical stability of the gravastar configuration.
5.2 Energy conditions
@@
As discussed earlier, in the braneworld scenario, the modifications to the gravitational field equations can be interpreted as effective contributions to the matter sector. Consequently, the standard energy conditions must be evaluated with respect to the effective stress-energy tensor rather than the original matter source. In this framework, the energy conditions are expressed as follows:
  • Null Energy Condition (NEC): $\rho^{\rm eff}\geq 0,$
  • Weak Energy Condition (WEC): $\rho^{\rm eff}+p_{r}^{\rm eff}\geq 0,\quad\rho^{\rm eff}+p_{t}^{\rm eff}\geq 0,$
  • Strong Energy Condition (SEC): $\rho^{\rm eff}+p_{r}^{\rm eff}+2p_{t}^{\rm eff}\geq 0,$
  • Dominant Energy Condition (DEC): $\rho^{\rm eff}-|p_{r}^{\rm eff}|\geq 0,\quad\rho^{\rm eff}-|p_{t}^{\rm eff}|\geq 0.$
@@
In Figure 4, we present the radial profiles of the various linear combinations of the effective energy density and pressures that correspond to these energy conditions across the finite shell. It is evident from the plots that all the physically relevant combinations remain strictly non-negative throughout the shell. This demonstrates that the effective matter supporting the gravastar satisfies all the standard energy conditions, ensuring both the physical plausibility and stability of the configuration.
@@
Figure 4: Variation of the different energy conditions w.r.t. the radial coordinate $r$.
6 Conclusions
@@
In this work, we have developed a gravastar model embedded in the Shtanov–Sahni braneworld, where the timelike nature of the extra dimension fundamentally alters the collapse dynamics. The SS scenario, with its negative brane tension and positive bulk cosmological constant, is distinguished by its ability to generate non-singular bouncing cosmologies. We have shown that this same mechanism, when applied to compact objects, provides a natural resolution of the singularity problem that plagues classical black hole solutions. Motivated by recent work on traversable wormholes in the SS braneworld [66], which showed that besides cosmological singularities, wormhole singularities can also be resolved in this framework, we demonstrate that gravastars too acquire a natural singularity-free description.
@@
Our gravastar construction exhibits several noteworthy features. The interior region is described by a Bose–Einstein condensate with equation of state $p=-\rho$, which contributes negatively to the active gravitational mass and dynamically provides the repulsive behavior required to prevent further gravitational collapse. In contrast to many earlier gravastar models that rely on idealized thin-shell approximations, the present framework admits exact analytic solutions for a shell of finite thickness composed of stiff matter, thereby offering a more realistic description of the transition layer. An additional intrinsic anisotropy arises from the projection of the bulk Weyl tensor onto the brane, leading to anisotropic effective pressures that emerge naturally within the Shiromizu–Maeda–Sasaki braneworld formalism and contribute to enhancing the overall stability of the configuration. Furthermore, the model consistently satisfies the required junction and energy conditions, and an analysis of the surface redshift confirms the physical viability and self-consistency of the resulting gravastar solution.
@@
The timelike extra dimension is central to these results. It ensures that the effective cosmological constant on the brane can vanish under fine-tuning, while simultaneously allowing the bulk corrections to regularize the interior geometry. This dual role highlights the SS braneworld as a particularly fertile ground for constructing singularity-free compact objects. Importantly, the induced anisotropy and suppressed effective mass are not ad hoc assumptions but direct consequences of the higher-dimensional corrections. Our findings resonate with recent investigations of traversable wormholes in the SS braneworld [66], where it was shown that wormhole curvature singularities can be avoided with matter satisfying the energy conditions. Taken together, these results suggest a unifying picture: the SS braneworld provides a consistent framework in which diverse singular structures in the form of cosmological bounces, wormholes, and gravastars can be rendered non-singular. This strengthens the case for the SS model as a promising candidate for probing quantum-gravity inspired phenomena in astrophysics.
@@
We briefly outline a few ideas concerning the possible observational signatures of gravastars, with particular emphasis on the braneworld framework. In scenarios where our four-dimensional universe is embedded in a higher-dimensional bulk spacetime, braneworld gravity has proven useful in interpreting several recent astrophysical observations. Notably, such models have been invoked in the context of the multimessenger event GW170817 observed by the LIGO/Virgo collaboration, along with its electromagnetic counterpart GRB170817A detected by the Fermi Gamma-ray Burst Monitor and the INTEGRAL Anti-Coincidence Shield spectrometer [67]. Furthermore, the observed shadow of the supermassive compact object M87 has been shown to admit a consistent explanation within a braneworld scenario featuring an Anti-de Sitter ($AdS_{5}$) bulk geometry [68]. These developments motivate the study of gravastars within frameworks that can be regarded as an ultraviolet-corrected extension of general relativity and that have already demonstrated relevance in describing non-singular black hole bounces [69] as well as non-singular bouncing cosmological models [34]. In this context, gravastars emerge as compelling horizonless alternatives to classical black holes, potentially compatible with quantum gravity–inspired modifications at high curvature scales.
@@
Although no direct observational evidence for gravastars currently exists, several indirect detection strategies have been proposed in the literature. These include the analysis of gravastar shadows [70], gravitational microlensing signatures exhibiting a higher maximal luminosity than those produced by black holes of identical mass [71], and characteristic features in gravitational-wave ringdown signals. In particular, it has been suggested that the ringdown phase of the GW150914 event [72] could, in principle, be generated by a horizonless compact object such as a gravastar [73, 74], although this interpretation remains under active debate [75]. Taken together, these studies indicate that while gravastars remain observationally elusive, current and future high-precision gravitational-wave and electromagnetic observations may provide viable avenues for distinguishing them from classical black holes, especially within modified gravity frameworks such as the braneworld scenario considered in the present work.
@@
In conclusion, the gravastar on the Shtanov–Sahni brane offers a compelling, physically consistent alternative to black holes, rooted in a framework that naturally resolves singularities. Our analysis reveals, in a fully analytic and self-consistent manner, how non-local bulk gravitational effects encoded through the projected Weyl tensor can dynamically generate intrinsic pressure anisotropy inside the gravastar, without the need for ad hoc anisotropic matter sources. This anisotropy, together with the repulsive contribution of the Bose–Einstein condensate core characterized by $p=-\rho$, plays a central gravitational role by weakening the effective attractive pull and enhancing the stability of the configuration against collapse. Unlike previous gravastar models, where stability is often imposed through phenomenological assumptions or fine-tuned junction conditions, here it emerges naturally from the modified gravitational sector itself, providing a genuinely geometric origin for the stabilizing mechanism. The existence of an exact finite-thickness stiff-matter shell further distinguishes our construction from past works relying on infinitesimally thin shells, allowing a more realistic description of intermediate regions and clarifying the physical interpretation of surface stresses, redshift bounds, and energy conditions in a higher-dimensional setting.

### v2

@@ V.1 Surface Redshift
$Z_{\mathrm{s}}$ $=$ $-1+\frac{1}{\sqrt{g_{\it tt}}}=-1+{\frac{1}{\sqrt{{C}^{2}{{e}^{Br^{2}}}}}}.$ (52)
@@
The radial profile of the surface redshift $Z_{\mathrm{s}}$ from the inner to the outer boundary of the gravastar is illustrated in Fig. 4. For any self-gravitating compact object, it is well known that the surface redshift must satisfy $Z_{\mathrm{s}}\leq 2$ Buchdahl1959 ; Straumann1984 ; Bohmer2006 in the absence of a cosmological constant. In our analysis within the Shtanov–Sahni braneworld framework, we do not include a $\Lambda$ term, as the effective four-dimensional cosmological constant vanishes due to a fine-tuned cancellation between the negative brane tension and the positive cosmological constant of the five-dimensional de Sitter bulk ($dS_{5}$).
@@
Using the numerically determined model parameters obtained from the matching and boundary conditions, the surface redshift reaches its maximum at the inner boundary of the gravastar and decreases smoothly across the shell, attaining a finite nonzero value at the outer boundary. The monotonic decrease of $Z_{\mathrm{s}}$ throughout the shell indicates a well-behaved gravitational potential, while its magnitude remains comfortably below the theoretical upper bound. This behavior not only confirms the regularity of the shell but also provides strong evidence for the physical stability of the gravastar configuration.
V.2 Energy conditions
@@
As discussed earlier, in the braneworld scenario, the modifications to the gravitational field equations can be interpreted as effective contributions to the matter sector. Consequently, the standard energy conditions must be evaluated with respect to the effective stress-energy tensor rather than the original matter source. In this framework, the energy conditions are expressed in Table II.
Figure 4: Surface redshift of the gravastar, $Z_{\mathrm{s}}$, plotted as a function of the radial coordinate $r$ across the shell region. The curves represents the numerical evaluation of Eq. (52) for different masses of the gravastar and for the parameter set $B=0.011415~{\rm km^{-2}}$ together with the constants determined from the junction conditions. The surface redshift decreases smoothly from the inner boundary of the shell toward the exterior boundary and remains well below the theoretical upper bound $Z_{\mathrm{s}}\leq 2$ for compact objects, indicating that the gravastar configuration is physically acceptable and gravitationally stable.
Table 2: Energy conditions.
Energy Condition Mathematical Form
Null Energy Condition (NEC) $\rho^{\mathrm{eff}}\geq 0$
Weak Energy Condition (WEC) $\rho^{\mathrm{eff}}+p_{r}^{\mathrm{eff}}\geq 0,\ \rho^{\mathrm{eff}}+p_{t}^{\mathrm{eff}}\geq 0$
Strong Energy Condition (SEC) $\rho^{\mathrm{eff}}+p_{r}^{\mathrm{eff}}+2p_{t}^{\mathrm{eff}}\geq 0$
Dominant Energy Condition (DEC) $\rho^{\mathrm{eff}}-|p_{r}^{\mathrm{eff}}|\geq 0,\ \rho^{\mathrm{eff}}-|p_{t}^{\mathrm{eff}}|\geq 0$
Table II summarizes the standard energy conditions used to assess the physical viability of the effective matter distribution in the gravastar model. Since braneworld corrections modify the gravitational field equations, these conditions are expressed in terms of the effective energy density $\rho_{\mathrm{eff}}$ and effective pressures $p_{r}^{\mathrm{eff}}$ and $p_{t}^{\mathrm{eff}}$. The table lists the mathematical forms of the null, weak, strong, and dominant energy conditions used in the analysis.
@@
In Fig. 5, we present the radial profiles of the various linear combinations of the effective energy density and pressures that correspond to these energy conditions across the finite shell. It is evident from the plots that all the physically relevant combinations remain strictly non-negative throughout the shell. This demonstrates that the effective matter supporting the gravastar satisfies all the standard energy conditions, ensuring both the physical plausibility and stability of the configuration.
@@
Figure 5: Radial variation of the combinations of effective energy density and pressures that define the standard energy conditions as functions of the radial coordinate $r$ within the gravastar shell. The plots are evaluated using the parameter values $B=0.011415~{\rm km^{-2}}$, $Q=0.138221$ and $\rho_{0}=0.05\times 10^{-7}~{\rm m}^{4}$, together with the constants obtained from the matching conditions. Here Fig. 5(A) [top-left panel] for $M=2.50M_{\odot}$, 5(B) [top-right panel]for $M=2.75M_{\odot}$, 5(C) [bottom-left panel]for $M=3.00M_{\odot}$, 5(D) [bottom-right panel] for $M=3.25M_{\odot}$.
VI Conclusions
@@
In this work, we have developed a gravastar model embedded in the SS braneworld, where the timelike nature of the extra dimension fundamentally alters the collapse dynamics. The SS scenario, with its negative brane tension and positive bulk cosmological constant, is distinguished by its ability to generate non-singular bouncing cosmologies. We have shown that this same mechanism, when applied to compact objects, provides a natural resolution of the singularity problem that plagues classical black hole solutions. Motivated by recent work on traversable wormholes in the SS braneworld Sengupta2023 , which showed that besides cosmological singularities, wormhole singularities can also be resolved in this framework, we demonstrate that gravastars too acquire a natural singularity-free description.
@@
Our gravastar construction exhibits several noteworthy features. The interior region is described by a Bose–Einstein condensate with equation of state $p=-\rho$, which contributes negatively to the active gravitational mass and dynamically provides the repulsive behavior required to prevent further gravitational collapse. In contrast to many earlier gravastar models that rely on idealized thin-shell approximations, the present framework admits exact analytic solutions for a shell of finite thickness composed of stiff matter, thereby offering a more realistic description of the transition layer. An additional intrinsic anisotropy arises from the projection of the bulk Weyl tensor onto the brane, leading to anisotropic effective pressures that emerge naturally within the Shiromizu–Maeda–Sasaki braneworld formalism SMS and contribute to enhancing the overall stability of the configuration. Furthermore, the model consistently satisfies the required junction and energy conditions, and an analysis of the surface redshift confirms the physical viability and self-consistency of the resulting gravastar solution.
@@
The timelike extra dimension is central to these results. It ensures that the effective cosmological constant on the brane can vanish under fine-tuning, while simultaneously allowing the bulk corrections to regularize the interior geometry. This dual role highlights the SS braneworld as a particularly fertile ground for constructing singularity-free compact objects. Importantly, the induced anisotropy and suppressed effective mass are not ad hoc assumptions but direct consequences of the higher-dimensional corrections. Our findings resonate with recent investigations of traversable wormholes in the SS braneworld Sengupta2023 , where it was shown that wormhole curvature singularities can be avoided with matter satisfying the energy conditions. Taken together, these results suggest a unifying picture: the SS braneworld provides a consistent framework in which diverse singular structures in the form of cosmological bounces, wormholes, and gravastars can be rendered non-singular. This strengthens the case for the SS model as a promising candidate for probing quantum-gravity inspired phenomena in astrophysics.
@@
We briefly outline a few ideas concerning the possible observational signatures of gravastars, with particular emphasis on the braneworld framework. In scenarios where our four-dimensional universe is embedded in a higher-dimensional bulk spacetime, braneworld gravity has proven useful in interpreting several recent astrophysical observations. Notably, such models have been invoked in the context of the multimessenger event GW170817 observed by the LIGO/Virgo collaboration, along with its electromagnetic counterpart GRB170817A detected by the Fermi Gamma-ray Burst Monitor and the INTEGRAL Anti-Coincidence Shield spectrometer Visinelli . Furthermore, the observed shadow of the supermassive compact object M87 has been shown to admit a consistent explanation within a braneworld scenario featuring an Anti-de Sitter ($AdS_{5}$) bulk geometry Vagnozzi . These developments motivate the study of gravastars within frameworks that can be regarded as an ultraviolet-corrected extension of general relativity and that have already demonstrated relevance in describing non-singular black hole bounces lqc as well as non-singular bouncing cosmological models Sahni4 . In this context, gravastars emerge as compelling horizonless alternatives to classical black holes, potentially compatible with quantum gravity–inspired modifications at high curvature scales.
@@
Although no direct observational evidence for gravastars currently exists, several indirect detection strategies have been proposed in the literature. These include the analysis of gravastar shadows Sakai2014 , gravitational microlensing signatures exhibiting a higher maximal luminosity than those produced by black holes of identical mass Kubo2016 , and characteristic features in gravitational-wave ringdown signals. In particular, it has been suggested that the ringdown phase of the GW150914 event Abbott2016 could, in principle, be generated by a horizonless compact object such as a gravastar Cardoso1 ; Cardoso2 , although this interpretation remains under active debate Chirenti2016 . Taken together, these studies indicate that while gravastars remain observationally elusive, current and future high-precision gravitational-wave and electromagnetic observations may provide viable avenues for distinguishing them from classical black holes, especially within modified gravity frameworks such as the braneworld scenario considered in the present work.
@@
In conclusion, the gravastar on the SS brane offers a compelling, physically consistent alternative to black holes, rooted in a framework that naturally resolves singularities. Our analysis reveals, in a fully analytic and self-consistent manner, how non-local bulk gravitational effects encoded through the projected Weyl tensor can dynamically generate intrinsic pressure anisotropy inside the gravastar, without the need for ad hoc anisotropic matter sources. This anisotropy, together with the repulsive contribution of the Bose–Einstein condensate core characterized by $p=-\rho$, plays a central gravitational role by weakening the effective attractive pull and enhancing the stability of the configuration against collapse. Unlike previous gravastar models, where stability is often imposed through phenomenological assumptions or fine-tuned junction conditions, here it emerges naturally from the modified gravitational sector itself, providing a genuinely geometric origin for the stabilizing mechanism. The existence of an exact finite-thickness stiff-matter shell further distinguishes our construction from past works relying on infinitesimally thin shells, allowing a more realistic description of intermediate regions and clarifying the physical interpretation of surface stresses, redshift bounds, and energy conditions in a higher-dimensional setting.